# Filtrated Grouping in Multiple Functional Regression

### v2

@@ 2.3 Filtrated Functional Partial Least Squares (filt-fPLS)
To identify the homogeneous components and estimate the coefficient scores, we developed a filtrated functional partial least squares method. To facilitate understanding of the proposed filt-fPLS algorithm, we begin with a simple illustrative case. In the ordinary grouped model (2-3), if $X_{1n}(t),\ldots,X_{pn}(t)\in\mathcal{K}_{1}$ (i.e., all covariates belong to a single group $\mathcal{K}_{1}$), then we can estimate the model using the classical functional partial least squares approach (see Delaigle and Hall [9]), where $\sum_{j\in\mathcal{K}_{1}}X_{jn}(t)$ is treated as the covariate. The first step is to maximize $\mbox{cov}(y_{n},\sum_{j\in\mathcal{K}_{1}}\langle X_{jn},\psi\rangle)$ over $\psi(t)$, subject to the constraint $\|\psi\|=1$. This is followed by a projection step, and the two steps are then repeated iteratively. In the filtrated grouped model (2-4), however, the grouping structure varies across dimensions. As a result, the maximization and projection steps are iteratively performed under different grouping configurations. The core principle of filt-fPLS is that the extracted homogeneous components should exhibit strong explanatory power for the response variable. We detail the algorithmic procedure below.
@@
$y_{n}^{[d]}$ $=y_{n}^{[d-1]}-\sum_{i=1}^{m_{d-1}}\zeta_{n,d-1}^{(i)}\hat{a}_{d-1,i},$
Particularly $X_{jn}^{[1]}(t)=X_{jn}(t)$, $y_{n}^{[1]}=y_{n}$.
It can be shown that under the forest structure, the homogeneous components across different layers are orthogonal (i.e., $\sum_{n\geq 1}\zeta_{nd}^{(i)}\zeta_{nd^{\prime}}^{(i^{\prime})}=0$ for $d\neq d^{\prime}$ and $i,i^{\prime}\geq 1$). In practical applications, the unknown covariances and expectations are replaced by their empirical counterparts, as discussed in Algorithm 1.
@@
2:while the ending criterion (discussed in Section 3.3) is not met do
@@
$\hat{\phi}_{d,i}(t)$ $=\arg\min_{\{\delta(t)\colon i\geq 1\}}\sum_{n=1}^{N}\left\|X_{jn}^{[d]}(t)-\sum_{i=1}^{m_{d}}\zeta_{nd}^{(i)}\delta_{i}(t)\right\|^{2},$
@@
3 Forest Structure Construction
The construction of the forest structure is based on the grouping path $\mathcal{P}$ derived from the ordinary grouping pursuit problem. Specifically, this involves identifying $\mathcal{K}_{1},\ldots,\mathcal{K}_{m}$ in model (2-3) by solving a regularized optimization problem over a sequence of increasing penalty parameters (see e.g., Shen and Huang [36]), and $\mathcal{P}$ encompasses all distinct grouping structures identified across different penalty parameters. Based on the grouping path, we first find a sequence of candidate grouping structures from which the forest structure is constructed. For the candidate grouping structures, it is required that the groups in each layer are nested within those of the preceding layer, as any non-nested grouping structures would complicate the construction of the forest structure. One challenge is that the grouping structures along the grouping path may not be nested as $\lambda$ increases, due to the occurrence of group splitting. Thus, it is crucial to carefully select the candidate grouping structures from the grouping path.
In the following, we discuss the identification of grouping path in Section 3.1, the selection of candidate grouping structures in Section 3.2, and the construction of the forest structure in Section 3.3.
3.1 Grouping Path
@@
where $\bm{\beta}(t)=\{\beta_{j}(t)\colon 1\leq j\leq p\}$, and $J_{\lambda}(\cdot)$ is some concave penalty, such as the SCAD penalty (see Fan and Li [12]) and the MCP (see Zhang [40]). As $\lambda$ increases, covariates tend to be clustered into fewer groups, although sometimes grouping splitting may occur. The sequence of grouping structures that emerge as the regularization parameter $\lambda$ increases defines the grouping path $\mathcal{P}$.
Suppose that for some orthonormal basis functions $\{\nu_{d}(t)\colon d\geq 1\}$, the following basis representations hold: $X_{jn}(t)=\sum_{d\geq 1}\xi_{jn,d}\nu_{d}(t)$ and $\beta_{j}(t)=\sum_{d\geq 1}b_{jd}\nu_{d}(t)$. Since $S(\bm{\beta}(t),\lambda)$ involves elements of infinite dimensions, we minimize the following truncated version over $\{b_{jd}\colon d=1,\ldots,D,\ j\geq 1\}$ with some sufficiently large $D\geq 1$,
$S_{D}(\bm{\beta}(t),\lambda)=\frac{1}{2}\sum_{n=1}^{N}\left(y_{n}-\sum_{j=1}^{p}\bm{\xi}^{\top}_{jn}\bm{B}_{j}\right)^{2}+\sum\limits_{i<j}J_{\lambda}(\|\bm{M}_{ij}\|),$ (3-5)
where $\bm{\xi}_{jn}=(\xi_{jn,1},\ldots,\xi_{jn,D})^{\top}$, $\bm{B}_{j}=(b_{j1},\ldots,b_{jD})^{\top}$ and $\bm{M}_{ij}=\bm{B}_{i}-\bm{B}_{j}$. We use the ADMM algorithm to solve (3-5). Let $\bm{B}=\{\bm{B}_{j}\colon j=1,\ldots,p\}$, $\bm{M}=\{\bm{M}_{ij}\colon i<j\}$, and
$H_{1}(\bm{B})=\frac{1}{2}\sum_{n=1}^{N}\left(y_{n}-\sum_{j=1}^{p}\bm{\xi}^{\top}_{jn}\bm{B}_{j}\right)^{2},\qquad H_{2}(\bm{M})=\sum_{i<j}J_{\lambda}(\|\bm{M}_{ij}\|),$
then the unconstrained optimization problem is equivalent to the following constrained optimization problem
$H(\bm{B},\bm{M})=H_{1}(\bm{B})+H_{2}(\bm{M}),\ \mbox{subject to}\ \bm{M}_{ij}=\bm{B}_{i}-\bm{B}_{j},\ 1\leq i<j\leq p.$ (3-6)
By the ADMM algorithm, the regularized estimates of $\bm{B},\bm{M}$ can be obtained by minimizing the augmented Lagrangian function
$L_{\theta}(\bm{B},\bm{M},\bm{u})$ $=H(\bm{B},\bm{M})+\sum_{i<j}\bm{u}^{\top}_{ij}(\bm{B}_{i}-\bm{B}_{j}-\bm{M}_{ij})+\frac{\theta}{2}\sum_{i<j}\|\bm{B}_{i}-\bm{B}_{j}-\bm{M}_{ij}\|^{2}$
over $\bm{B}$ and $\bm{M}$. Given $\bm{B}$ and multipliers $\bm{u}=\{u_{ij,dd^{\prime}},i<j,d<d^{\prime}\}$, we update $\bm{M}$ by minimizing $L_{\theta}(\bm{B},\bm{M},\bm{u})$ over $\bm{M}$, and the optimization problem is equivalent to minimizing the following function over $\bm{M}$,
$\frac{\theta}{2}\sum_{i<j}\|\bm{M}_{ij}-(\bm{B}_{i}-\bm{B}_{j})-\theta^{-1}\bm{u}_{ij}\|^{2}+\sum_{i<j}J_{\lambda}(\|\bm{M}_{ij}\|).$ (3-7)
For $j=1,\ldots,p$, define $\bm{\Xi}_{j}=\left(\bm{\xi}_{1j},\ldots,\bm{\xi}_{N}\right)^{\top}$, $\bm{\Xi}=(\bm{\Xi}_{1}|\ldots|\bm{\Xi}_{p})$. The minimizer of $\bm{B}$ given $\bm{M}$ and $\bm{u}$ of the augmented Lagrangian function is
$\bm{B}=(\bm{\Xi}^{\top}\bm{\Xi}+\theta\bm{\Delta}^{\top}\bm{\Delta})^{-1}(\bm{\Xi}^{\top}\bm{y}+\theta\bm{\Delta}^{\top}(\bm{M}-\theta^{-1}\bm{u})).$ (3-8)
Here, $\bm{\Delta}=(\bm{\Delta}_{ij},i<j)$ represents the matrix formed by concatenating the row blocks $\bm{\Delta}_{ij}$, and $\bm{\Delta}_{ij}=(\cdots\bm{0},\bm{I},\bm{0},\cdots,\bm{0},-\bm{I},\bm{0}\cdots)$ is a $d\times pd$ matrix where $\bm{I}$ is the $d\times d$ identity matrix. We summarize the above discussion in Algorithm 2.
Algorithm 2 The ADMM algorithm for grouping pursuit
1:Initialize estimates $\bm{B}^{(0)}$ (e.g., ordinary least squares estimate of $\bm{B}$) and set $\bm{u}^{(0)}=\bm{0}$.
2:while convergence criterion is not met do
3:     Given $\bm{B}^{(m)}$ and $\bm{u}^{(m)}$, calculate $\bm{M}^{(m+1)}$ by minimizing (3-7) over $\bm{M}$.
4:     Given $\bm{M}^{(m+1)}$ and $\bm{u}^{(m)}$, calculate $\bm{B}^{(m+1)}$ using (3-8).
5:     Update the multipliers $u^{(m+1)}_{ij,dd^{\prime}}=u^{(m)}_{ij,dd^{\prime}}+\theta\{b^{(m+1)}_{id}-b^{(m+1)}_{jd}-M^{(m+1)}_{ij,d}\}$.
6:end while
7:return $\bm{B}^{(m+1)}$.
3.2 Candidate Grouping Structures
@@
After obtaining the grouping path $\mathcal{P}$, we first construct candidate sets of grouping structures from the path, from which the forest structure is constructed. Notationally, for two grouping structures $G,G^{\prime}$, $G\subseteq G^{\prime}$ means that every group in $G$ is nested in some group in $G^{\prime}$. Denote a candidate set by $\mathcal{G}=\{G_{1},\ldots,G_{\ell}\}$, where $\ell$ is the number of grouping structures in $\mathcal{G}$, and $G_{i}\in\mathcal{P}$ denotes the $i$-th layer’s grouping structure of the candidate set $\mathcal{G}$ satisfying $G_{\ell}\subseteq G_{\ell-1}\subseteq\cdots\subseteq G_{1}$.
@@
Multiple candidate sets can be constructed from the grouping path, and only one of them is used for constructing the forest structure of model (2-4). The relationship $\mathcal{G}\subseteq\mathcal{G}^{\prime}$ means that every grouping structure in $\mathcal{G}$ is also included in $\mathcal{G}^{\prime}$. To make the candidate set as inclusive as possible, we require that any two different candidate sets $\mathcal{G}=\{G_{1},\ldots,G_{\ell}\}$ and $\mathcal{G}^{\prime}=\{G^{\prime}_{1},\ldots,G^{\prime}_{\ell^{\prime}}\}$ satisfy that a) $\mathcal{G}\not\subseteq\mathcal{G}^{\prime}$ and $\mathcal{G}^{\prime}\not\subseteq\mathcal{G}$, b) ${G}_{1}={G}^{\prime}_{1}=\{1,2,\ldots,p\}$, and ${G}_{\ell}={G}^{\prime}_{\ell^{\prime}}=\{\{1\},\{2\},\ldots,\{p\}\}$. The selected candidate set should ensure sufficient diversity in grouping structures to support forest construction and achieve strong predictive performance in the resulting model. We now present the procedure for selecting the candidate set.
@@
    Final Selection: Compute the average prediction error of the models built across all grouping structures within each candidate set and select the one that yields the lowest average prediction error.
@@
This selection procedure ensures that the chosen candidate set represents a comprehensive forest hierarchy, capturing the finest level of details available in the grouping process. Meanwhile, the grouping structures in the selected candidate set retain strong explanatory power, thereby aligning with the covariance maximization principle central to filt-fPLS.
3.3 Iterative Selection of Grouping Structures
The forest structure is constructed from the selected candidate set $\widetilde{\mathcal{G}}=\{\widetilde{G}_{1},\ldots,\widetilde{G}_{\ell}\}$ through an iterative process. Specifically, starting from the first layer, we sequentially examine all the grouping structure in the selected candidate set $\widetilde{G}_{1},\ldots,\widetilde{G}_{\ell}$. If, based on a predefined criterion, the grouping structure selected for the current layer is $\widetilde{G}_{i}$, the process continues to the next layer by sequentially testing $\widetilde{G}_{i},\ldots,\widetilde{G}_{\ell}$.
@@ 3.3 Iterative Selection of Grouping Structures
We develop a generalized information criterion (GIC, see e.g., Nishii [31] and Zhang et al. [41]) to select the grouping structures across different layers. The format of GIC criterion is: GIC = measure of model fit $+$ tuning parameter $\times$ measure of model complexity. In principle, if the covariates are separated into too many groups, the resulting grouping structure would fail to explain the covariate interdependence. Therefore, we penalize complex grouping structure. Notationally, denote a forest structure by $\mathcal{F}$, and $F_{d}\in\widetilde{\mathcal{G}}$ as the $d$-th layer’s grouping structure in $\mathcal{F}$. We develop an iterative GIC criterion, with the layer-wise GIC value defined as
$\mbox{GIC}(F_{d})=N^{-1}\sum_{n=1}^{N}(y^{[d+1]}_{n})^{2}+\tau_{d}|F_{d}|,$
@@
where $|F_{d}|$ denotes the number of groups in $F_{d}$, and $y^{[d+1]}_{n}$’s are the ($d+1$)-th layer’s response residuals in the filt-fPLS procedure based on the grouping structure $F_{d}$. We select $F_{d}$ to minimize $\mbox{GIC}(F_{d})$. Here, $\{\tau_{d}\colon d\geq 1\}$ is a non-increasing sequence with respect to $d$, thus the number of groups tends to increase as $d$ increases. Clearly, denote $\widetilde{\mathcal{F}}$ as the identified forest structure, we have the relationship ~$\widetilde{\mathcal{F}}\subseteq\widetilde{\mathcal{G}}\subseteq\mathcal{P}$.
The number of tuning parameters $\{\tau_{d}\colon d\geq 1\}$ increases with $d$, making tuning parameter selection increasingly complicated when the total number of layers is large. To achieve efficient tuning parameter adaptation, we propose adopting a parametric form for $\tau_{d}$, for example, $\tau_{\theta}(d)=\rho\gamma^{-d+1}$, where $\theta=(\rho,\gamma)$. Using a parametric form significantly reduces the computational burden, as it requires selecting only $\theta$ rather than individually determining each $\tau_{d}$.
@@
Ending criterion. Now we discuss how to fix the maximal layer of the forest structure. In principle, the filtration should continue until the extracted homogeneous components no longer significantly improve the response’s fitting error. Thus, we propose the following ending criterion: we stop filtration at the $d$-th layer if $\sum_{n}\{(y_{n}^{[d]})^{2}-(y_{n}^{[d+1]})^{2}\}/\sum_{n}(y_{n}^{[d]})^{2}<e$, where $e>0$ is a small value. If $\sum_{n}\|X^{[d+1]}_{jn}\|^{2}$ becomes sufficiently small, we remove the $j$-th covariate from further filtration.
4 Simulation

### v3

@@ 2.3 Filtrated Functional Partial Least Squares (filt-fPLS)
To extract the homogeneous components and estimate the coefficient scores, we develope a filtrated functional partial least squares method. To facilitate understanding of the proposed filt-fPLS algorithm, we begin with a simple illustrative case. In the ordinary grouped model (2-3), if $X_{1n}(t),\ldots,X_{pn}(t)\in\mathcal{K}_{1}$ (i.e., all covariates belong to a single group $\mathcal{K}_{1}$), then we can estimate the model using the classical functional partial least squares approach (see Delaigle and Hall [9]), where $\sum_{j\in\mathcal{K}_{1}}X_{jn}(t)$ is treated as the covariate. The first step is to maximize $\mbox{cov}(y_{n},\sum_{j\in\mathcal{K}_{1}}\langle X_{jn},\psi\rangle)$ over $\psi(t)$, subject to the constraint $\|\psi\|=1$. The resulting feature $\sum_{j\in\mathcal{K}_{1}}\langle X_{jn},\psi\rangle$ is used to explain the response. This is followed by a projection step, and the two steps are then repeated iteratively. In the filtrated grouped model (2-4), however, the grouping structure varies across dimensions. As a result, the maximization and projection steps are iteratively performed under different grouping configurations. The core principle of filt-fPLS is that the extracted homogeneous components should exhibit strong explanatory power for the response variable. We detail the algorithmic procedure below.
@@
$y_{n}^{[d]}$ $=y_{n}^{[d-1]}-\sum_{i=1}^{m_{d-1}}\zeta_{n,d-1}^{(i)}\hat{a}_{d-1,i}.$
Particularly, $X_{jn}^{[1]}(t)=X_{jn}(t)$, $y_{n}^{[1]}=y_{n}$.
In each layer, the filt‑fPLS basis functions and the corresponding coefficient scores jointly account for homogeneous association patterns of covariates in the same group. Note that under forest structure, the homogeneous components across different layers are orthogonal (i.e., $\sum_{n\geq 1}\zeta_{nd}^{(i)}\zeta_{nd^{\prime}}^{(i^{\prime})}=0$ for $d\neq d^{\prime}$ and $i,i^{\prime}\geq 1$). Consequently, the homogeneous components identified at different layers capture distinct information, allowing the filtrated grouped model (2-4) to represent multi-resolution kinematic connectivity structure without redundancy and with clearer interpretability. In practical applications, the unknown covariances and expectations are replaced by their empirical counterparts, as discussed in Algorithm 1. In the next section, we discuss how to identify the forest structure.
@@
2:while the ending criterion (discussed in Section 2.4.3) is not met do
@@
$\hat{\phi}_{d,i}(t)$ $=\arg\min_{\{\delta_{i}(t)\colon i\geq 1\}}\sum_{n=1}^{N}\left\|X_{jn}^{[d]}(t)-\sum_{i=1}^{m_{d}}\zeta_{nd}^{(i)}\delta_{i}(t)\right\|^{2},$
@@
2.4 Forest Structure Construction
The construction of the forest structure is based on the grouping path $\mathcal{P}$ derived from the ordinary grouping pursuit problem. Specifically, this involves identifying $\mathcal{K}_{1},\ldots,\mathcal{K}_{m}$ in model (2-3) by solving a regularized optimization problem over a sequence of increasing penalty parameters (see e.g., Shen and Huang [34]), and $\mathcal{P}$ encompasses all distinct grouping structures identified across different penalty parameters. We apply filtration along the grouping path because covariates that share strong homogeneity in their coefficient functions tend to merge earlier, providing useful information on the hierarchical structures of kinematic connectivity. Based on the grouping path, we first find a sequence of candidate grouping structures from which the forest structure is constructed. For the candidate grouping structures, it is required that the groups in each layer are nested within those of the preceding layer, as any non-nested grouping structures would complicate the construction of the forest structure. One challenge is that the grouping structures along the grouping path may not be nested as $\lambda$ increases, due to the occurrence of group splitting. Thus, it is crucial to carefully select the candidate grouping structures from the grouping path. In the following, we discuss the identification of grouping path in Section 2.4.1, the selection of candidate grouping structures in Section 2.4.2, and the construction of the forest structure in Section 2.4.3.
2.4.1 Grouping Path
@@
where $\bm{\beta}(t)=\{\beta_{j}(t)\colon j=1,\ldots,p\}$, and $J_{\lambda}(\cdot)$ is some concave penalty, such as the SCAD penalty (see Fan and Li [12]) and the MCP (see Zhang [39]). As $\lambda$ increases, covariates tend to be clustered into fewer groups, although occasional group splits may still occur. The sequence of all grouping structures that emerge as the regularization parameter $\lambda$ increases defines the grouping path $\mathcal{P}$. Suppose that for some orthonormal basis functions $\{\nu_{d}(t)\colon d\geq 1\}$, the following basis representations hold $X_{jn}(t)=\sum_{d\geq 1}\xi_{jn,d}\nu_{d}(t)$ and $\beta_{j}(t)=\sum_{d\geq 1}b_{jd}\nu_{d}(t)$. Since $S(\bm{\beta}(t),\lambda)$ involves infinite-dimensional elements, we minimize the following truncated version over $\{b_{jd}\colon d=1,\ldots,D,\ j\geq 1\}$ with some sufficiently large $D\geq 1$,
$S_{D}(\bm{\beta}(t),\lambda)=\frac{1}{2}\sum_{n=1}^{N}\left(y_{n}-\sum_{j=1}^{p}\bm{\xi}^{\top}_{jn}\bm{B}_{j}\right)^{2}+\sum\limits_{i<j}J_{\lambda}(\|\bm{B}_{i}-\bm{B}_{j}\|),$ (2-5)
where $\bm{\xi}_{jn}=(\xi_{jn,1},\ldots,\xi_{jn,D})^{\top}$, $\bm{B}_{j}=(b_{j1},\ldots,b_{jD})^{\top}$. We use the ADMM algorithm to minimize (2-5). The details are included in the supplementary materials.
2.4.2 Candidate Grouping Structures
@@
After obtaining the grouping path $\mathcal{P}$, we first construct candidate sets of grouping structures from the path, from which the forest structure is then constructed. Notationally, for two grouping structures $^{\prime}, GGG\subseteq G^{\prime}$ means that every group in $G$ is nested in some group in $G^{\prime}$. Denote a candidate set by $\mathcal{G}=\{G_{1},\ldots,G_{\ell}\}$, where $\ell$ is the number of grouping structures in $\mathcal{G}$, and $G_{i}\in\mathcal{P}$ denotes the $i$-th layer’s grouping structure of the candidate set $\mathcal{G}$ satisfying $G_{\ell}\subseteq G_{\ell-1}\subseteq\cdots\subseteq G_{1}$.
@@
Multiple candidate sets can be constructed from the grouping path, and only one of them is used for constructing the forest structure of model (2-4). The relationship $\mathcal{G}\subseteqq\mathcal{G}^{\prime}$ means that every grouping structure in $\mathcal{G}$ is also included in $\mathcal{G}^{\prime}$. To make the candidate set as inclusive as possible, we require that any two different candidate sets $\mathcal{G}=\{G_{1},\ldots,G_{\ell}\}$ and $\mathcal{G}^{\prime}=\{G^{\prime}_{1},\ldots,G^{\prime}_{\ell^{\prime}}\}$ satisfy that a) $\mathcal{G}\not\subseteqq\mathcal{G}^{\prime}$ and $\mathcal{G}^{\prime}\not\subseteqq\mathcal{G}$, b) ${G}_{1}={G}^{\prime}_{1}=\{1,2,\ldots,p\}$, and ${G}_{\ell}={G}^{\prime}_{\ell^{\prime}}=\{\{1\},\{2\},\ldots,\{p\}\}$. The selected candidate set should ensure sufficient diversity in grouping structures to support forest construction and achieve strong predictive performance in the resulting model. We now present the procedure for selecting the candidate set.
@@
    Final Selection: For each candidate set, compute the average prediction error of the models built across all the included grouping structures, and select the set that yields the lowest average prediction error.
@@
This selection procedure ensures that the chosen candidate set represents a comprehensive forest hierarchy, including both the finest and the coarsest grouping. Meanwhile, the grouping structures in the selected candidate set retain strong explanatory power, thereby aligning with the covariance maximization principle central to filt-fPLS.
2.4.3 Iterative Selection of Grouping Structures
@@
The forest structure is constructed from the selected candidate set $\widetilde{\mathcal{G}}=\{\widetilde{G}_{1},\ldots,\widetilde{G}_{\ell}\}$ through an iterative process. If the grouping structure selected for the current layer is $\widetilde{G}_{i}$ based on a predefined criterion, the process continues to the next layer by sequentially testing $\widetilde{G}_{i},\ldots,\widetilde{G}_{\ell}$. Particularly, we sequentially examine all the grouping structures in $\widetilde{\mathcal{G}}$ in the first layer of filtration. Regarding the predefined criterion, we adopt the generalized information criterion (GIC, see e.g., Zhang et al. [40]) to select the grouping structures across different layers. The format of GIC criterion is: GIC = measure of model fit ++ tuning parameter $\times$ measure of model complexity. In principle, if the covariates are separated into too many groups, the resulting grouping structure would fail to explain the kinematic connectivity. Therefore, we penalize complex grouping structure.
Denote a forest structure by $\mathcal{F}$, and $F_{d}\in\widetilde{\mathcal{G}}$ as the $d$-th layer’s grouping structure in $\mathcal{F}$. We develop an iterative GIC criterion, with the layer-wise GIC value defined as
$\mbox{GIC}(F_{d})=N^{-1}\sum\limits_{n=1}^{N}(y^{[d+1]}_{n})^{2}+\tau_{d}|F_{d}|,$
@@
where $|F_{d}|$ denotes the number of groups in $F_{d}$, and $y^{[d+1]}_{n}$’s are the ($d+1$)-th layer’s response residuals in the filt-fPLS procedure based on the grouping structure $F_{d}$. We select $F_{d}$ to minimize $\mbox{GIC}(F_{d})$. Here, $\{\tau_{d}\colon d\geq 1\}$ is a non-increasing sequence with respect to $d$, thus the number of groups tends to increase as $d$ increases. The number of tuning parameters $\{\tau_{d}\colon d\geq 1\}$ increases with $d$, making tuning parameter selection increasingly complicated when the total number of layers is large. To solve this issue, we propose adopting a parametric form for $\tau_{d}$ such as $\tau_{\theta}(d)=\rho\gamma^{-d+1}$, where $\theta=(\rho,\gamma)$. The parametric form significantly reduces the computational burden, as it requires selecting only $\theta$ rather than individually determining each $\tau_{d}$. Clearly, denote $\widetilde{\mathcal{F}}$ as the identified forest structure, we have the relationship $\widetilde{\mathcal{F}}\subseteqq\widetilde{\mathcal{G}}\subseteqq\mathcal{P}$.
@@
We now turn to the stopping criterion for the filtrated grouping procedure. In principle, the filtration should continue until the extracted homogeneous components no longer significantly improve the model’s explanatory power. Thus, we propose the following ending criterion: we stop filtration at the $d$-th layer if $\sum_{n}\{(y_{n}^{[d]})^{2}-(y_{n}^{[d+1]})^{2}\}/\sum_{n}(y_{n}^{[d]})^{2}<e$, where $e>0$ is a small value. In addition, if $\sum_{n}\|X^{[d+1]}_{jn}\|^{2}$ becomes sufficiently small, the $j$-th covariate is removed from filtration starting at the $(d+1)$-th layer to reduce computational burden.
3 Simulation results